OpenStudy (anonymous):

write sin x + sin 5x as a product

jimthompson5910 (jim_thompson5910):

Hint: Use the identity \[\large \sin(u) + \sin(v) = 2\sin\left(\frac{u+v}{2}\right)\cos\left(\frac{u-v}{2}\right)\]

OpenStudy (anonymous):

thanks!

jimthompson5910 (jim_thompson5910):

you're welcome, tell me what you get

OpenStudy (anonymous):

idk if i'm doing this right! sin (1) + sin (5) =2sin3cos-3

jimthompson5910 (jim_thompson5910):

\[\large \sin(u) + \sin(v) = 2\sin\left(\frac{u+v}{2}\right)\cos\left(\frac{u-v}{2}\right)\] \[\large \sin(x) + \sin(5x) = 2\sin\left(\frac{x+5x}{2}\right)\cos\left(\frac{x - 5x}{2}\right)\] \[\large \sin(x) + \sin(5x) = 2\sin\left(\frac{6x}{2}\right)\cos\left(\frac{-4x}{2}\right)\] \[\large \sin(x) + \sin(5x) = 2\sin\left(3x\right)\cos\left(-2x\right)\] \[\large \sin(x) + \sin(5x) = 2\sin\left(3x\right)\cos\left(2x\right)\]

jimthompson5910 (jim_thompson5910):

btw I found that identity here http://www.sosmath.com/trig/Trig5/trig5/trig5.html

OpenStudy (anonymous):

thanksssss

jimthompson5910 (jim_thompson5910):

sure thing

OpenStudy (anonymous):

Now that i'm looking at it, it wasn't difficult at all! The only difficult part was knowing what formula to use.

jimthompson5910 (jim_thompson5910):

yeah it's not too bad

OpenStudy (anonymous):

would you be willing to help me solve csc (cos^-1 (15/17))

jimthompson5910 (jim_thompson5910):

draw a right triangle and mark one of the angles (that's not a right angle) as \(\large \theta\) 

jimthompson5910 (jim_thompson5910):

now cos^-1(15/17) refers to the fact that there is some angle \(\large \theta\) such that \(\large \cos(\theta) = \frac{15}{17}\)

jimthompson5910 (jim_thompson5910):

since cos = adj/hyp, we know that 

jimthompson5910 (jim_thompson5910):

how would you find the opposite side